Batteries use a variety of metals. Based on the Lab Introduction and the table near the...

  1. Batteries use a variety of metals. Based on the Lab Introduction and the table near the end of the lab and determine whether zinc and manganese are close in activity or far apart. Discuss a reason for using different metals.

  2. Develop a grid needed to test each of the metals in Part I with each of the solutions. Sketch this in your lab notebook. Explain how you can systematically test the metals without running duplicate trials.

  3. Sodium metal in water is a popular demonstration. Watch one of the YouTube videos of this demonstration. Explain 2 safety concerns with having sodium metal in the laboratory. Describe 2 indications that a chemical reaction is occurring.

5. Use the table of standard reduction potentials in your text, or another suitable reference, to complete the following table. Write the first reaction as an oxidation and the second as a reduction. An example is provided.

Electrodes

Half-reactions

Zn Cu

Zn(s) → Zn2+(aq) + 2e– Cu2+(aq) + 2e– → Cu(s)

+0.76 V +0.34 V

Cu Pb

Pb Ag

Pb Mg

Pb Zn

6. The batteries we commonly use incorporate heavy metals. Identify the heavy metals that must be disposed of in the electrode waste. Choose one of the heavy metals and explain the effect of the heavy metal on the environment.

SHOW WORK PLEASE ,thank you !
Answer #1

Answer 4 Please provide metals being used to form a grid.

Safety concerns of using sodium:

a) It should be kept under kerosene or paraffin as it can catch fire on exposure to air.

b) It should not be touched with bare hands as it can react causing burns.

Indications for chemical reaction occurring:

a) Bubbling in the reaction mixture due to evolution of gas.

b) Dissolution of sodium metal in the solution .

Answer 5

Electrodes

Half-reactions

Zn Cu

Zn(s) → Zn2+(aq) + 2e–,

Cu2+(aq) + 2e– → Cu(s)

+0.76 V +0.34 V =+ 1.10 V

Cu Pb

Pb(s) → Pb2+(aq) + 2e–

Cu2+(aq) + 2e– → Cu(s)

+0.13 V + 0.34 V = +0.47 V

Pb Ag

Pb(s) → Pb2+(aq) + 2e–

2Ag+(aq) + 2e– → 2Ag(s)

+0.13 V + 0.80 V = +0.93 V

Pb Mg

Mg(s) → Mg2+(aq) + 2e–

Pb2+(aq) + 2e– → Pb(s)

+2.37 V - 0.13 V = +2.24 V

Pb Zn

Zn(s) → Zn2+(aq) + 2e–

Pb2+(aq) + 2e– → Pb(s)

+0.76 V - 0.13 V = +0.63V

Answer 6

One of the heavy metal that is mostly used is Pb (Lead)

1) Lead remains as dust forever in the environment thus entering through respiration cycle.

2) Dissolved lead contaminates water bodies and hence enter food chains with aquatic organisms.

3) Entry of lead dust in soil contaminates it and leads to toxication of crop products.